1. Arrange s, p and d sub-shells of a shell in the increasing order of effective nuclear charge (Zeff) experienced by the electron present in them. 

2. Nickel atom can lose two electrons to form Ni2+ ion. The atomic number of nickel is 28. From which orbital will nickel lose two electrons.

3. Calculate the total number of angular nodes and radial nodes present in 3porbital.



1
Expert's answer
2020-07-22T07:59:22-0400

1. order for effective nuclear charge : s>p>d because s is closest to nucleus so it has maximum attraction from nucleus while d is far from nucleus so it feel less attraction from nucleus.


2 Atomic number of Ni is 28 and its electronic configuration is "1s^22s^22p^63s^23p^63d^84s^2."

So Ni will lose electron from 4s orbital.


3. no. of radial nodes = n-l-1

for 3p; n= 3 and l = 1 for p-orbital

So no. of radial nodes = 3-1-1 = 1

no. of angular nodes = l = 1